Solve the initial value problem below using the method of Laplace transforms. 2ty - 5ty + 5y = 20, y(0) = 4, y0) = -3 Clic
Answer #1

24 y - Y51-20 ५८०) =५.५०) = -3 ++ - व ( c s? YCS) – 400 =960)) 1 yen - 4543) ds -1 -३१) +4 -SEYCO) 1 (3) ।। _ त (()) I (SYDate._// 8-१० (-252+5:) Ycs) (45 - 10) Y) = १-8 S(25-5) y(g) + 225-5)YC) = नकम 2 8.5-20 +17) = 5 (२:-5) = ५ (23-2 SC25-5) न